DEVELOPMENT OF BROAD RANGE SCAN CAPABILITIES WITH JET COOLED CAVITY RINGDOWN SPECTROSCOPY

\underline{TERRANCE J. CODD}, MING-WEI CHEN and TERRY A. MILLER, Laser Spectroscopy Facility, The Ohio State University, Columbus, Ohio 43210.

We have developed a technique for obtaining broad scans, >100 cm^{−1}, for jet cooled cavity ringdown spectroscopy (CRDS) spectra. Previously the scans of the jet cooled, CRDS apparatus were limited to <10 cm^{−1} due to the use of a narrow linewidth radiation source. However, by coupling our jet cooled, CRDS apparatus with a moderate resolution (≃ 0.05 cm^{−1}) dye laser we are able to greatly increase our rate of data acquisition thereby gaining the capability to perform broad spectral surveys of jet cooled molecules. As a test of the capabilities of the technique we have scanned the Ã–X̃ transition of NO_{3} previously reported by Deev et al.^{a} at room temperature. We believe that this will be a very useful technique to search for transitions of cold molecules whose frequencies are not well known and which later can be studied using high resolution methods.

aA. Deev, J. Sommar, and M. Okumura, J. Chem. Phys. 122, 224305 (2005).